-} e-2t 2 [

solve it with 3 different ways.
first, use method of undetermined coefficient
second, use diagonalization
third, use laplace transform
'please solve this problem as precise as possible!'
Answer #1

-D -2t e. 3 3xy -6e2t Coefficieut 3 t2 -6 e DA3)x 6 e2t x(D+3) -p43)213 O+3)222. 6e =-He Ths s (D)Fe), hre f (D) = DsD+g in dm mt 2 m +8 -0 m(m++) +2(m+H)-0, m=21- けて . PE) Pauticulor Iwtegsal e (a) 4e2 -2t (D+(D+) -H E2+H) (DA2) -2 te26 eneral Solut